# A Theory of the NEPv Approach for Optimization On the Stiefel Manifold

### v2

@@ 5 Convex Composition
$T_{3}\,:\,P\in\mathbb{O}^{n\times k}\to T_{3}(P):=\begin{bmatrix}\operatorname{tr}((P_{1}^{\operatorname{T}}A_{1}P_{1})^{m_{1}})\\ \vdots\\ \operatorname{tr}((P_{\ell}^{\operatorname{T}}A_{\ell}P_{\ell})^{m_{\ell}})\\ \operatorname{tr}((P_{\ell+1}^{\operatorname{T}}D_{1})^{m_{\ell+1}})\\ \vdots\\ \operatorname{tr}((P_{\ell+t}^{\operatorname{T}}D_{t})^{m_{\ell+t}})\\ \end{bmatrix}\in\mathbb{R}^{\ell+t},$ (5.11)
where integer $m_{i}\geq 1$ for all $1\leq i\leq\ell+t$. It reduces to Example 5.1 if $m_{i}=1$ for all $1\leq i\leq\ell+t$. Either $\ell=0$ or $t=0$ is allowed. Suppose all $A_{i}\succeq 0$. If all $P_{i}$ together have the properties in (5.7), then the consistency conditions in (5.5) are satisfied with
=$\mathbb{P}=\{P\in\mathbb{O}^{n\times k}\,:\,P_{\ell+j}^{\operatorname{T}}D_{j}\succeq 0\,\,\mbox{for $1\leq j\leq t$}\},$ (5.12)
$\alpha=1$, $\beta_{i}=2m_{i}-1$ for $1\leq i\leq\ell$ and $\beta_{\ell+j}=m_{\ell+j}-1$ for $1\leq j\leq t$, by Theorems 4.3 and 4.4. Theorem 5.1 applies, assuming $\phi_{j}(\boldsymbol{x})\geq 0$ for $1\leq j\leq\ell$ and for $\ell+1\leq j\leq\ell+t$ such that $m_{j}\geq 2$.
@@
$T_{1a}\,:\,P\in\mathbb{O}^{n\times k}\to T_{1a}(P):=\begin{bmatrix}\operatorname{tr}(P^{\operatorname{T}}A_{1}P)\\ \vdots\\ \operatorname{tr}(P^{\operatorname{T}}A_{\ell}P)\\ \operatorname{tr}(P^{\operatorname{T}}D_{1})\\ \vdots\\ \operatorname{tr}(P^{\operatorname{T}}D_{t})\\ \end{bmatrix}\in\mathbb{R}^{\ell+t}.$ (5.13)
@@
Consider $f=\phi\circ T_{1a}$ where $\phi$ is convex and differentiable with $\phi_{j}(\boldsymbol{x})\geq 0$ for $1\leq j\leq\ell$, and suppose $A_{i}\succeq 0$ for $1\leq i\leq\ell$. Given $\widehat{P}\in\mathbb{O}^{n\times k},\,P\in\mathbb{O}^{n\times k}$, let $\widetilde{P}=\widehat{P}Q$ where $Q$ is an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{D}(P)$, where
@@
$T_{1}$ and $T_{1a}$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}=I_{k}$. $T_{1}$ with (5.7) and $t\geq 1$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}$ is implicitly determined by $\widetilde{P}_{\ell+j}^{(i+1)}=\widehat{P}_{\ell+j}^{(i)}S_{j}$ where $S_{j}$ is an orthogonal polar factor of $\phi_{\ell+j}(T_{1}(P))(\widehat{P}_{\ell+j}^{(i)})^{\operatorname{T}}D_{j}$ for $1\leq j\leq t$. $T_{1}$ with (5.8) and $t\geq 1$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}$ is implicitly determined by: for just one element $j$ from each $\mathbb{I}_{q}$ ($1\leq q\leq\tau$), $\widetilde{P}_{\ell+j}^{(i+1)}=\widehat{P}_{\ell+j}^{(i)}S_{q}$ where $S_{q}$ is an orthogonal polar factor of $\left[\widehat{P}_{\ell+p}^{(i)}\right]^{\operatorname{T}}\left[\sum_{p\in\mathbb{I}_{q}}\phi_{\ell+p}(T_{1}(P))D_{p}\right]$ for $1\leq q\leq\tau$. $T_{1a}$ with $t\geq 1$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}$ is an orthogonal polar factor of $(\widehat{P}^{(i)})^{\operatorname{T}}\mathscr{D}(P^{(i)})$. $T_{2}$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell+t$, $Q_{i}=I_{k}$. $T_{3}$ with $t=0$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}=I_{k}$. $T_{3}$ with (5.7) and $t\geq 1$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$ and for $\ell+1\leq j\leq\ell+t$ such that $m_{j}\geq 2$, $Q_{i}$ is implicitly determined by $\widetilde{P}_{\ell+j}^{(i+1)}=\widehat{P}_{\ell+j}^{(i)}S_{j}$ where $S_{j}$ is an orthogonal polar factor of $(\widehat{P}_{\ell+j}^{(i)})^{\operatorname{T}}D_{j}$ for $1\leq j\leq t$.
@@
In Table 3, we list conditions on partial derivatives $\phi_{j}$ and the best choices of $Q_{i}$ at Line 4 of Algorithm 1 that, when it is not $I_{k}$, can increase the objective value even more per SCF iterative step than the NPDo Ansatz suggests. Having said that, we notice that $P^{(i)}$ as $i$ varies may belong to different subsets $\mathbb{P}$ of $\mathbb{O}^{n\times k}$. For example, with $T=T_{1a}$, if $Q_{i}$ is calculated according to Theorem 5.2, i.e., $Q_{i}$ is an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}\mathscr{D}(P^{(i)})$, then $P^{(i+1)}\in\mathbb{P}_{i}:=\{X\in\mathbb{O}^{n\times k}\,:\,X^{\operatorname{T}}\mathscr{D}(P^{(i)})\succeq 0\}$ that varies from one iterative step to the next. Eventually, $\mathbb{P}_{i}$ approaches :=$\mathbb{P}_{*}:=\{X\in\mathbb{O}^{n\times k}\,:\,X^{\operatorname{T}}\mathscr{D}(P_{*})\succeq 0\}$ by Corollary 5.1. Similar comments can be said about $T_{1}$ with (5.7) or (5.8) that we discussed towards the end of Example 5.1. Numerically, such variations in $\mathbb{P}$ does not pose any problem for Algorithm 1 to compute an approximate maximizer for the maximization problem (5.1).
@@
We conclude this section by commenting on the applicability of the results of this section to the objective functions in Table 1 via convex compositions of atomic functions for NPDo. Essentially our results are applicable to all of those that are not in the quotient form, i.e., SEP, MBSub, SumCT, TrCP, UMDS, and DFT, assuming that matrices $A$ and $A_{i}$ are positive semidefinite. SEP is simply about the atomic function $\operatorname{tr}(P^{\operatorname{T}}AP)$. For OLDA and SumTR, the corresponding composing functions $\phi$ are $x_{2}/x_{1}$ and $x_{2}/x_{1}+x_{3}$, respectively, but both are non-convex. The composing function for OCCA is $\phi_{0}(\boldsymbol{x})=x_{2}/\sqrt{x_{1}}$ where $\boldsymbol{x}=[x_{1},x_{2}]^{\operatorname{T}}$, which is not convex but whose square $\phi(\boldsymbol{x}):=[\phi_{0}(\boldsymbol{x})]^{2}=x_{2}^{2}/x_{1}$ is convex for $x_{2}\geq 0$ and $x_{1}>0$. Unfortunately, the atomic function associated with $x_{1}$ is $\operatorname{tr}(P^{\operatorname{T}}BP)$, for which $\phi_{1}(\boldsymbol{x}):=\partial\phi(\boldsymbol{x})/\partial x_{1}=-(x_{2}/x_{1})^{2}\leq 0$, violating the conditions of Theorems 5.1 and 5.2. A similar argument applies to $\Theta$TR.
@@ 6 The NEPv Framework
for $P\in\mathbb{O}^{n\times k}$ (or even $\mathbb{R}^{n\times k}$), where $H(P)\in\mathbb{R}^{n\times n}$ is a symmetric matrix-valued function dependent of $P$, e.g., for $f(P)=\frac{\operatorname{tr}(P^{\operatorname{T}}AP)}{\operatorname{tr}(P^{\operatorname{T}}BP)}+\operatorname{tr}(P^{\operatorname{T}}CP)$ from the sum of trace ratios (SumTR) [75, 76], which includes SEP and OLDA as special cases,
@@
In the case when $\mathscr{H}(P)\equiv H_{0}(P)P$, this $H(P)$ becomes $2\,H_{0}(P)$ for $P\in\mathbb{O}^{n\times k}$.
@@
Let $H(P)\in\mathbb{R}^{n\times n}$ be a symmetric matrix-valued function on $\mathbb{O}^{n\times k}$, satisfying
$H(P)P-\frac{\partial f(P)}{\partial P}=P\,\mathscr{M}(P)\quad\mbox{for $P\in\mathbb{O}^{n\times k}$},$ (6.6)
where $\mathscr{M}(P)\in\mathbb{R}^{k\times k}$ is some matrix-valued function. $P\in\mathbb{O}^{n\times k}$ is a solution to the KKT condition (2.3) if and only if it is a solution to NEPv (6.3) and $\mathscr{M}(P)$ is symmetric.
@@
If $P$ is a solution to the KKT condition (2.3), i.e., $\mathscr{H}(P)=P\Lambda$, $P\in\mathbb{O}^{n\times k}$, and $\Lambda=\Lambda^{\operatorname{T}}$. Then, by (6.6),
@@ 6.1 The NEPv Ansatz
The successes of the NEPv approach used in [67, 77] for solving OCCA and $\Theta$TR relies on certain monotonicity lemmas which inspire us to make the following ansatz to build our framework upon. It also requires a sufficiently inclusive subset $\mathbb{P}$ of $\mathbb{O}^{n\times k}$ as in the NPDo framework.
The NEPv Ansatz. For function $f$ defined in some neighborhood $\mathbb{O}_{\delta}^{n\times k}$ of $\mathbb{O}^{n\times k}$, there is a symmetric matrix-valued function $H(P)\in\mathbb{R}^{n\times n}$ such that for $\widehat{P}\in\mathbb{O}^{n\times k},\,P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$, if
@@
then there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+\omega\eta$, where $\omega$ is some positive constant, independent of $P$ and $\widehat{P}$.
For any $P\in\mathbb{O}^{n\times k}$, there is always $\widehat{P}\in\mathbb{O}^{n\times k}$ such that (6.7) holds with some $\eta>0$, unless for that given $P$, $H(P)P=P\Omega$ holds and the eigenvalues of $\Omega$ consist of the $k$ largest eigenvalues of $H(P)$. In fact, we can take $\widehat{P}\in\mathbb{O}^{n\times k}$ to be an orthonormal basis matrix of the eigenspace of $H(P)$ associated with its $k$ largest eigenvalues, which also maximizes $\operatorname{tr}(X^{\operatorname{T}}H(P)X)$ over $X\in\mathbb{O}^{n\times k}$. In general, it is the desirable aim, $f(\widetilde{P})\geq f(P)+\omega\eta$, in the NEPv Ansatz that needs to be verified before the general theory of this section can be applied. Below we will use the same objective function in Example 3.1 to rationalize this ansatz.
@@
$H(P)P-\mathscr{H}(P)=P\big{[}2(D^{\operatorname{T}}P)^{2}\big{]})\quad\mbox{for}\quad P\in\mathbb{O}^{n\times k},$
satisfying (6.6). To achieve equivalency between the KKT condition (2.3) and NEPv (6.3), according to Theorem 6.1 we should limit the scope to those $P\in\mathbb{O}^{n\times k}$ such that $(D^{\operatorname{T}}P)^{2}$ is symmetric. Actually we will further limit the scope to $P\in\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$. Suppose now that (6.7) holds for $P\in\mathbb{O}_{D+}^{n\times k}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, or equivalently,
@@
Next let $Q\in\mathbb{O}^{k\times k}$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$ and let $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$. We find that $\operatorname{tr}(\widetilde{P}^{\operatorname{T}}A\widetilde{P})=\operatorname{tr}(\widehat{P}^{\operatorname{T}}A\widehat{P})$, but it remains to break the second term in the left-hand side of (6.8) apart so that $P$ and $\widehat{P}$ are detached. For that purpose, we note
@@
$\leq 2\|\widehat{P}^{\operatorname{T}}DP^{\operatorname{T}}D\|_{\operatorname{tr}}$
@@
    The NEPv Ansatz critically involves a symmetric matrix-valued function $H(P)\in\mathbb{R}^{n\times n}$ that has to be constructed. In Theorem 6.1, we provide a guiding equation (6.6) for the purpose so that the KKT condition (2.3) and NEPv (6.3) are equivalent as far as solving the associated optimization problem (1.1) is concerned. It is fulfilled naturally when $\mathscr{H}(P)\equiv H(P)P$ for $P\in\mathbb{O}^{n\times k}$ exactly (e.g., for OLDA, SumTR, TrCP, UMDS, and DFT), but at other times, we will have to construct $H(P)$ individually based on the particularity of $\mathscr{H}$ as in [77] for OCCA, [67] for $\Theta$TR, [74] for MAXBET and Example 6.1, or we simply use the generic (6.5). The ansatz does not demand any explicit association of $H(P)$ with $\mathscr{H}(P)$, but conceivably they should be highly related, such as the relation imposed by (6.6).
@@
    This is a stronger version, however, assuming that for any $\widehat{P}\in\mathbb{O}^{n\times k}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$. Here is why. Suppose that (6.10) holds. Given $\widehat{P}\in\mathbb{O}^{n\times k},\,P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$, let $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ for some $Q\in\mathbb{O}^{k\times k}$. If (6.7) holds, then
@@
    When $f(P)$ is right-unitarily invariant, $H(P)$ always exists such that (6.1) holds and can be taken to be right-unitarily invariant, too [46, Lemma 2.1]. In such a case, the ansatz can be simplified to: $Q=I_{k}$ and $\widetilde{P}=\widehat{P}$ always because $f(\widehat{P})=f(\widetilde{P})$ regardless of $Q$. Also often $\mathbb{P}=\mathbb{O}^{n\times k}$.
@@
    Introducing a subset $\mathbb{P}$ of $\mathbb{O}^{n\times k}$ and judiciously choosing $Q$ are for generality to deal with the case when $f(P)$ is not right-unitarily invariant, e.g., the one in Example 6.1 and those in Table 1 in section 1 that involve $D$ or $D_{i}$. Suitable $Q$ can also increase the objective value more than $\omega\eta$. For example, for $\Theta$TR with $H(P)$ given by (6.4), the NEPv Ansatz holds with taking $Q$ to be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$. In fact, along the line of the proof of [67, Lemma 2.1], assuming $B\succeq 0$, $\operatorname{rank}(B)>n-k$, and $\operatorname{tr}(P^{\operatorname{T}}AP+P^{\operatorname{T}}D)\geq 0$ in the case of $10<\theta<$ 1 but otherwise no condition on $\operatorname{tr}(P^{\operatorname{T}}AP+P^{\operatorname{T}}D)$ for $\theta\in\{0,1\}$, we can improve the conclusion of [67, Theorem 2.2] to (in the current notation)

### v4

@@ 5 Convex Composition
$T_{3}\,:\,P\in{\rm St}(k,n)\to T_{3}(P):=\begin{bmatrix}\operatorname{tr}((P_{1}^{\operatorname{T}}A_{1}P_{1})^{m_{1}})\\ \vdots\\ \operatorname{tr}((P_{\ell}^{\operatorname{T}}A_{\ell}P_{\ell})^{m_{\ell}})\\ \operatorname{tr}((P_{\ell+1}^{\operatorname{T}}D_{1})^{m_{\ell+1}})\\ \vdots\\ \operatorname{tr}((P_{\ell+t}^{\operatorname{T}}D_{t})^{m_{\ell+t}})\\ \end{bmatrix}\in\mathbb{R}^{\ell+t},$ (5.11)
where integer $m_{i}\geq 1$ for all $1\leq i\leq\ell+t$. It reduces to Example 5.1 if $m_{i}=1$ for all $1\leq i\leq\ell+t$. Either $\ell=0$ or $t=0$ is allowed. Suppose all $A_{i}\succeq 0$. Suppose all $P_{i}$ together have the properties in (5.7). Then the consistency conditions in (5.5) are satisfied with
=$\mathbb{P}=\{P\in{\rm St}(k,n)\,:\,P_{\ell+j}^{\operatorname{T}}D_{j}\succeq 0\,\,\mbox{for $1\leq j\leq t$}\},$ (5.12)
$\alpha=1$, $\beta_{i}=2m_{i}-1$ for $1\leq i\leq\ell$ and $\beta_{\ell+j}=m_{\ell+j}-1$ for $1\leq j\leq t$, by Theorems 4.3 and 4.4. Theorem 5.1 applies, assuming $\phi_{j}(\boldsymbol{x})\geq 0$ for $1\leq j\leq\ell$ and for each $j\in\{\ell+1,\ldots,\ell+t\}$ with $m_{j}\geq 2$. $Q$ is implicitly determined by $\widetilde{P}_{\ell+j}=\widehat{P}_{\ell+j}S_{j}$ where $S_{j}$ is an orthogonal polar factor of $\widehat{P}_{\ell+j}^{\operatorname{T}}D_{j}$ for $1\leq j\leq t$.
@@
$T_{1a}\,:\,P\in{\rm St}(k,n)\to T_{1a}(P):=\begin{bmatrix}\operatorname{tr}(P^{\operatorname{T}}A_{1}P)\\ \vdots\\ \operatorname{tr}(P^{\operatorname{T}}A_{\ell}P)\\ \operatorname{tr}(P^{\operatorname{T}}D_{1})\\ \vdots\\ \operatorname{tr}(P^{\operatorname{T}}D_{t})\\ \end{bmatrix}\in\mathbb{R}^{\ell+t}.$ (5.13)
@@
Consider $f=\phi\circ T_{1a}$ where $\phi$ is convex and differentiable with $\phi_{j}(\boldsymbol{x})\geq 0$ for $1\leq j\leq\ell$, and suppose $A_{i}\succeq 0$ for $1\leq i\leq\ell$. Given $\widehat{P}\in{\rm St}(k,n),\,P\in{\rm St}(k,n)$, let $\widetilde{P}=\widehat{P}Q$ where $Q$ is an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{D}(P)$ with
@@
$T_{1}$ and $T_{1a}$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}=I_{k}$. $T_{1}$ with (5.7) and $t\geq 1$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}$ is implicitly determined by $\widetilde{P}_{\ell+j}^{(i+1)}=\widehat{P}_{\ell+j}^{(i)}S_{j}$ where $S_{j}$ is an orthogonal polar factor of $\phi_{\ell+j}(T_{1}(P))\big{[}\widehat{P}_{\ell+j}^{(i)}\big{]}^{\operatorname{T}}D_{j}$ for $1\leq j\leq t$. $T_{1}$ with (5.8) and $t\geq 1$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}$ is implicitly determined by: for just one element $j$ from each $\mathbb{I}_{q}$ ($1\leq q\leq\tau$), $\widetilde{P}_{\ell+j}^{(i+1)}=\widehat{P}_{\ell+j}^{(i)}S_{q}$ where $S_{q}$ is an orthogonal polar factor of $\big{[}\widehat{P}_{\ell+j}^{(i)}\big{]}^{\operatorname{T}}\big{[}\sum_{p\in\mathbb{I}_{q}}\phi_{\ell+p}(T_{1}(P))D_{p}\big{]}$ for $1\leq q\leq\tau$. $T_{1a}$ with $t\geq 1$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}$ is an orthogonal polar factor of $\big{[}\widehat{P}^{(i)}\big{]}^{\operatorname{T}}\mathscr{D}(P^{(i)})$. $T_{2}$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell+t$, $Q_{i}=I_{k}$. $T_{3}$ with $t=0$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$, $Q_{i}=I_{k}$. $T_{3}$ with (5.7) and $t\geq 1$ $\phi_{j}\geq 0$ for $1\leq j\leq\ell$ and for each $j\in\{\ell+1,\ldots,\ell+t\}$ with $m_{j}\geq 2$, $Q_{i}$ is implicitly determined by $\widetilde{P}_{\ell+j}^{(i+1)}=\widehat{P}_{\ell+j}^{(i)}S_{j}$ where $S_{j}$ is an orthogonal polar factor of $\big{[}\widehat{P}_{\ell+j}^{(i)}\big{]}^{\operatorname{T}}D_{j}$ for $1\leq j\leq t$. * $\phi_{j}(\boldsymbol{x}):=\partial\phi(\boldsymbol{x})/\partial x_{j}$ for $\boldsymbol{x}=[x_{j}]$.
@@
In Table 3, we list conditions on partial derivatives $\phi_{j}$ and the best choices of $Q_{i}$ at Line 4 of Algorithm 1 that, when it is not $I_{k}$, can increase the objective value even more per SCF iterative step than the NPDo Ansatz suggests. Having said that, we notice that $P^{(i)}$ as $i$ varies may belong to different subsets $\mathbb{P}$ of ${\rm St}(k,n)$. For example, with $T=T_{1a}$, if $Q_{i}$ is calculated according to Theorem 5.2, i.e., $Q_{i}$ is an orthogonal polar factor of $\big{[}\widehat{P}^{(i)}\big{]}^{\operatorname{T}}\mathscr{D}(P^{(i)})$, then $P^{(i+1)}\in\mathbb{P}_{i}:=\{X\in{\rm St}(k,n)\,:\,X^{\operatorname{T}}\mathscr{D}(P^{(i)})\succeq 0\}$ that varies from one iterative step to the next. Eventually, $\mathbb{P}_{i}$ approaches :=$\mathbb{P}_{*}:=\{X\in{\rm St}(k,n)\,:\,X^{\operatorname{T}}\mathscr{D}(P_{*})\succeq 0\}$ by Corollary 5.1. Similar comments can be said about $T_{1}$ with (5.7) or (5.8) that we discussed towards the end of Example 5.1. Numerically, such variations in $\mathbb{P}$ does not pose any problem for Algorithm 1 to compute an approximate maximizer for the maximization problem (5.1).
@@
We conclude this section by commenting on the applicability of the results of this section to the objective functions in Table 1 via convex compositions of atomic functions for NPDo. Essentially our results are applicable to all of those that are not in the quotient form, i.e., SEP, MBSub, SumCT, TrCP, UMDS, and DFT, assuming that matrices $A$ and $A_{i}$ are positive semidefinite. SEP is simply about the atomic function $\operatorname{tr}(P^{\operatorname{T}}AP)$. For OLDA and SumTR, the corresponding composing functions $\phi$ are $x_{2}/x_{1}$ and $x_{2}/x_{1}+x_{3}$, respectively, but both are non-convex. The composing function for OCCA is $\phi_{0}(\boldsymbol{x})=x_{2}/\sqrt{x_{1}}$ where $\boldsymbol{x}=[x_{1},x_{2}]^{\operatorname{T}}$, which is not convex but whose square $\phi(\boldsymbol{x}):=[\phi_{0}(\boldsymbol{x})]^{2}=x_{2}^{2}/x_{1}$ is convex for $x_{2}\geq 0$ and $x_{1}>0$. Unfortunately, the atomic function associated with $x_{1}$ is $\operatorname{tr}(P^{\operatorname{T}}BP)$, for which $\phi_{1}(\boldsymbol{x}):=\partial\phi(\boldsymbol{x})/\partial x_{1}=-(x_{2}/x_{1})^{2}\leq 0$, violating the conditions of Theorems 5.1 and 5.2. A similar argument applies to $\Theta$TR. So the NPDo approach does not work for OLDA, OCCA, and $\Theta$TR. But, fortunately, the NEPv approach next will.
@@ 6 The NEPv Framework
for $P\in{\rm St}(k,n)$ (or even $\mathbb{R}^{n\times k}$), where $H(P)\in\mathbb{R}^{n\times n}$ is a symmetric matrix-valued function dependent of $P$, e.g., for $f(P)=\frac{\operatorname{tr}(P^{\operatorname{T}}AP)}{\operatorname{tr}(P^{\operatorname{T}}BP)}+\operatorname{tr}(P^{\operatorname{T}}CP)$ from the sum of trace ratios (SumTR) [75, 76], which includes SEP and OLDA as special cases,
@@
In the case when $\mathscr{H}(P)\equiv H_{0}(P)P$, this $H(P)$ becomes $2\,H_{0}(P)$ for $P\in{\rm St}(k,n)$.
@@
Let $H(P)\in\mathbb{R}^{n\times n}$ be a symmetric matrix-valued function satisfying
$H(P)P-\frac{\partial f(P)}{\partial P}=P\,\mathscr{M}(P)\quad\mbox{for $P\in{\rm St}(k,n)$},$ (6.6)
where $\mathscr{M}(P)\in\mathbb{R}^{k\times k}$ is some matrix-valued function. $P\in{\rm St}(k,n)$ is a solution to the KKT condition (2.3) if and only if it is a solution to NEPv (6.3) and $\mathscr{M}(P)$ is symmetric.
@@
If $P$ is a solution to the KKT condition (2.3), i.e., $\mathscr{H}(P)=P\Lambda$, $P\in{\rm St}(k,n)$, and $\Lambda=\Lambda^{\operatorname{T}}$. Then, by (6.6),
@@ 6.1 The NEPv Ansatz
The successes of the NEPv approach used in [67, 77] for solving OCCA and $\Theta$TR relies on certain monotonicity lemmas which inspire us to make the following ansatz to build our framework upon. It also requires a sufficiently inclusive subset $\mathbb{P}$ of ${\rm St}(k,n)$ as in the NPDo framework in Part I.
The NEPv Ansatz. For function $f$ defined in some neighborhood ${\rm St}_{\delta}(k,n)$ of the Stiefel manifold ${\rm St}(k,n)$, there is a symmetric matrix-valued function $H(P)\in\mathbb{R}^{n\times n}$ such that for $\widehat{P}\in{\rm St}(k,n),\,P\in\mathbb{P}\subseteq{\rm St}(k,n)$, if
@@
then there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+\omega\eta$, where $\omega$ is some positive constant, independent of $P$ and $\widehat{P}$.
For any $P\in{\rm St}(k,n)$, there is always $\widehat{P}\in{\rm St}(k,n)$ such that (6.7) holds with some $\eta>0$, unless for that given $P$, $H(P)P=P\Omega$ holds and the eigenvalues of $\Omega$ consist of the $k$ largest eigenvalues of $H(P)$. In fact, we can take $\widehat{P}\in{\rm St}(k,n)$ to be an orthonormal basis matrix of the eigenspace of $H(P)$ associated with its $k$ largest eigenvalues, which also maximizes $\operatorname{tr}(X^{\operatorname{T}}H(P)X)$ over $X\in{\rm St}(k,n)$ [20]. Hence, for the purpose of solving (1.1), we may relax the ansatz to $\eta\geq 0$ only. In general, it is the desirable aim, $f(\widetilde{P})\geq f(P)+\omega\eta$, in the NEPv Ansatz that needs to be verified before the general theory of this section can be applied. Below we will use the same objective function in Example 3.1 to rationalize this ansatz.
@@
$H(P)P-\mathscr{H}(P)=P\big{[}2(D^{\operatorname{T}}P)^{2}\big{]})\quad\mbox{for}\quad P\in{\rm St}(k,n),$
satisfying (6.6). To achieve equivalency between the KKT condition (2.3) and NEPv (6.3), according to Theorem 6.1 we should limit the scope to those $P\in{\rm St}(k,n)$ such that $(D^{\operatorname{T}}P)^{2}$ is symmetric. Actually we will further limit the scope to $P\in\mathbb{P}={\rm St}(k,n)_{D+}$. Suppose now that (6.7) holds for $P\in{\rm St}(k,n)_{D+}$ and $\widehat{P}\in{\rm St}(k,n)$, or equivalently,
@@
Next let $Q\in{\rm St}(k,k)$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$ and let $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$. We find that $\operatorname{tr}(\widetilde{P}^{\operatorname{T}}A\widetilde{P})=\operatorname{tr}(\widehat{P}^{\operatorname{T}}A\widehat{P})$, but it remains to break the second term in the left-hand side of (6.8) apart so that $P$ and $\widehat{P}$ are detached. For that purpose, we note
@@
$\leq 2\|\widehat{P}^{\operatorname{T}}DP^{\operatorname{T}}D\|_{\operatorname{tr}}\qquad(\mbox{since $\|P^{\operatorname{T}}\widehat{P}\|_{2}\leq 1$})$
@@
    The NEPv Ansatz critically involves a symmetric matrix-valued function $H(P)\in\mathbb{R}^{n\times n}$ that has to be constructed. In Theorem 6.1, we provide a guiding equation (6.6) for the purpose so that the KKT condition (2.3) and NEPv (6.3) are equivalent as far as solving the associated optimization problem (1.1) is concerned. It is fulfilled naturally when $\mathscr{H}(P)\equiv H(P)P$ for $P\in{\rm St}(k,n)$ exactly (e.g., for OLDA, SumTR, TrCP, UMDS, and DFT), but at other times, we will have to construct $H(P)$ individually based on the particularity of $\mathscr{H}$ as in [77] for OCCA, [67] for $\Theta$TR, [74] for MAXBET and Example 6.1, or we simply use the generic (6.5). The ansatz does not demand any explicit association of $H(P)$ with $\mathscr{H}(P)$, but conceivably they should be highly related, such as the relation imposed by (6.6).
@@
    This is a stronger version, however, assuming that for any $\widehat{P}\in{\rm St}(k,n)$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$. Here is why. Suppose that (6.10) holds. Given $\widehat{P}\in{\rm St}(k,n),\,P\in\mathbb{P}\subseteq{\rm St}(k,n)$, let $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ for some $Q\in{\rm St}(k,k)$. If (6.7) holds, then
@@
    When $f(P)$ is right-unitarily invariant, $H(P)$ always exists such that (6.1) holds and can be taken to be right-unitarily invariant, too [46, Lemma 2.1]. In such a case, the ansatz can be simplified to: $Q=I_{k}$ and $\widetilde{P}=\widehat{P}$ always because $f(\widehat{P})=f(\widetilde{P})$ regardless of $Q$. Also often $\mathbb{P}={\rm St}(k,n)$.
@@
    Introducing a subset $\mathbb{P}$ of ${\rm St}(k,n)$ and judiciously choosing $Q$ are for generality to deal with the case when $f(P)$ is not right-unitarily invariant, e.g., the one in Example 6.1 and those in Table 1 in section 1 that involve $D$ or $D_{i}$. Suitable $Q$ can also increase the objective value more than $\omega\eta$. For example, for $\Theta$TR with $H(P)$ given by (6.4), the NEPv Ansatz holds with taking $Q$ to be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$. In fact, along the line of the proof of [67, Lemma 2.1], assuming $B\succeq 0$, $\operatorname{rank}(B)>n-k$, and $\operatorname{tr}(P^{\operatorname{T}}AP+P^{\operatorname{T}}D)\geq 0$ in the case of $10<\theta<$ 1 but otherwise no need to impose nonnegativity on $\operatorname{tr}(P^{\operatorname{T}}AP+P^{\operatorname{T}}D)$ for $\theta\in\{0,1\}$, we can improve the conclusion of [67, Theorem 2.2] to (in the current notation)